OpenStudy (anonymous):

Graph : Center: Vertices: Foci: Equation Below:

OpenStudy (anonymous):

\[\frac{(x+3)^2}{ 4 } + \frac{ (y-1)^2 }{ 25 } = 1\]

OpenStudy (anonymous):

Can anyone help?

hero (hero):

Did you lose your textbook? How come you're not able to do this?

OpenStudy (anonymous):

I missed this topic because the teacher went over it for 5 minutes and didn't mention it again.. && I need to have this section done by tonight. @Hero

OpenStudy (anonymous):

Can you guide me through it? @Peter14

OpenStudy (mertsj):

Which conic section is this?

OpenStudy (anonymous):

Ellipse

OpenStudy (mertsj):

Is the major axis horizontal or vertical?

OpenStudy (anonymous):

Vertical

OpenStudy (mertsj):

What is the center?

OpenStudy (mertsj):

\[\frac{(x-h)^2}{b^2}+\frac{(y-k)^2}{a^2}=1\]

OpenStudy (mertsj):

That is the equation of an ellipse whose center is (h,k) and whose major axis is vertical.

OpenStudy (mertsj):

What is the center of your ellipse?

OpenStudy (anonymous):

3, -1?

OpenStudy (mertsj):

Notice that the center is (h,k) and both h and k follow a negative sign.

OpenStudy (anonymous):

-3, -1? or would the 1 be positive because it's already negative?

OpenStudy (mertsj):